OpenStudy (anonymous):

3y-6/x=12 if y=4 for some reason i keep getting 144 but thats not right….

OpenStudy (anonymous):

(3(4)-6)/x = 12

OpenStudy (anonymous):

6/x = 12

OpenStudy (anonymous):

x = 1/2

OpenStudy (anonymous):

Assuming it's this equation.. \[\frac{ 3y-6 }{ x } = 12\]

OpenStudy (anonymous):



OpenStudy (anonymous):

its that lol

OpenStudy (anonymous):

What's the answer?

OpenStudy (anonymous):

answer choices are A. 0 B. -1/4 C. 6 D. -4

OpenStudy (anonymous):

i have no idea

OpenStudy (anonymous):